Solve the equation 7^(x – 2) = 5^x. Round the answer to four decimal places.
a. 8.3219
b. 9.6253
c. 10.1263
d. 11.5665
Thank you for your help.

Answer by stanbon(75887)

Solve the equation 7^(x – 2) = 5^x. Round the answer to four decimal places.
--------------
(x-2)log(7) = xlog(5)
---
xlog(7) - xlog(5) = 2log(7)
x[log(7)-log(5)] = 2log(7)
x = [2log(7)]/[log(7)-log(5)[
---
x = 11.5665
==================
Cheers,
Stan H.
=================
a. 8.3219
b. 9.6253
c. 10.1263
d. 11.5665